Shift Summary for Eve shift, 1/11/00


Subject: Shift Summary for Eve shift, 1/11/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:20:31 CST


              Shift Summary for Eve shift, 1/11/00

On Shift: ST, BobT
Also: EB, SGlazov, Arun, JasonL, RickK, and others

Hours of beam delivered at all intensities: 8h
Hours of beam delivered at nominal intensity: 8h
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 4h

Accounting of lost time:
  Lots of SHORT_REG832 startup problems:
    L3 modifications, monitoring fixes, readout problems,
    trigger modifications,..., you name it, we got it.

========================================================================
Runs ended this shift:
  Run 15480 - 366K events - BEAM_MUON_SHORT
  Run 15481 - 15K online w/cuts reg 2pi events - SHORT_REG832
  Run 15483 - 1.5K online w/cuts reg 2pi events - SHORT_REG832

In progress:
  Run 15484 - 50K online w/cuts reg 2pi events - SHORT_REG832

Tapes brought to FCC this shift:
 NONE

========================================================================
Accesses during shift:
  BobT and ST to remove HA latch & TDC crate from readout.

========================================================================
Overview of shift:
  Short regenerator running. ~66K online w/cuts reg 2pi events
  collected by end of shift.

========================================================================
Description of Shift:

------------------------------------------------------------------------

16:15 Start of first SHORT_REG832 run 15481 at 3E12.
       
       Various problems, notably:

         o Many DAQ_DEAD errors due to full memories
           and high average cpu time/event of ~7ms.
  
         o Monitoring of reg 2pi events screwed up.
                        
       H dibaryon trigger rates (corrected for PS and deadtime):
  
         L1: 9.75 KHz/E12
         L2: 7.42 KHz/E12

18:00 Stop run to make various changes:

         Fix reg 2pi monitoring.
         Remake L3 to speed up processing of H dibaryon triggers.
         Update H dibaryon trigger (see XXIII.230).

18:19 Start SHORT_REG832 run 15483 at 2.8E12.

19:00 Sextupole PS problems. Stop run.

       In an effort to reduce data volume,
       the HA latch and TDC crate (stream 5) and the
       bananas (stream 6) are removed from the readout.

       Further changes made to H dibaryon trigger (XXIII.232).

19:16 Start SHORT_REG832 run 15484.

       Having bananas out of readout causes triggers
       to hang -- put back in.

20:40 Crate 80 taken out of readout => no MA ADCs for now.

       Problems with KQ/BAN readout. JasonL comes in
       to remove X-view trig sources; trigger redefined as of
       spill 139. 3HC2X and 2HC2X removed from triggers 11 and 12.
       
       Data volume in stream 6 is now significantly less.

22:40 Sporadic stream 5 problems in DC-1Y TDC crate -- CTI word
       appears to be dropping "4"s bit. (FFF8 should be FFFC).

23:00 Intensity request reduced to 2.4E12.

       RickK continues work to reduce FERA readout deadtime.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

  Debug DC-1Y TDC crate CTI word problem during downtime.
  CTI word appears to be dropping "4"s bit. (FFF8 should be FFFC).



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:20:35 CST